OpenStudy (anonymous):

find the partial derivative f(x,y) = sin(x-y) at (-1,-1)! xD

OpenStudy (anonymous):

\(\frac{\partial f}{\partial x}\)? or \(\frac{\partial f}{\partial y}\)? $$f=\sin(x-y)\\f_x=\cos(x-y)\\f_y=-\cos(x-y)$$

OpenStudy (anonymous):

$$f_x(-1,-1)=\cos(0)=1\\ f_y(-1,-1)=-\cos(0)=-1\\ \nabla f=<1,-1>=i-j$$

OpenStudy (anonymous):

thank you so much!!!! :)